Question 3 0 out of 10 poin Consider an M/M/T system with Lambda of 9.68 and Mu ot 15.39 What is the probability that a customer will find a queue upon arrival at the system? Please show work
Answer #1

Probability of having no customer in the queue = (1-lambda/mue) = (1-9.68/ 5.99)=0.616

Probability of having customers in the queue =1-0.616 =0.384

So the answer is =0384